OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = −x2 + 5x + 4 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 3 to x = 5 represent?

OpenStudy (anonymous):

The water falls down with an average speed of 3 feet per second from 3 seconds to 5 seconds. The water falls down with an average speed of 5 feet per second from 3 seconds to 5 seconds. The water travels an average distance of 3 feet from 3 seconds to 5 seconds. The water travels an average distance of 5 feet from 3 seconds to 5 seconds.

OpenStudy (anonymous):

@campbell_st plzz

OpenStudy (anonymous):

@welshfella anything

OpenStudy (anonymous):

please bro

OpenStudy (welshfella):

plug in x = 3 and x = 5 into the equation to see the height of the water at these times f)3) = -(3)^2 + 5(3) + 4 = -9+15+4 = 10 feet can you work out the height for x = 5?

OpenStudy (welshfella):

just plug in x=5 and work it out as above

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

what exactly do i have to solve

OpenStudy (welshfella):

do as I have done but replace the 3 by 5 f(5) = -(5)^2 etc

OpenStudy (anonymous):

ughh

OpenStudy (welshfella):